OpenStudy (anonymous):

Crystal earns $5.25 per hour mowing lawns. Write a rule to describe how the amount of money m earned is a function of the number of hours h spent mowing lawns. How much does Crystal earn if she works 2 hours and 15 minutes?

OpenStudy (thomas5267):

\[ \text{Money earned}=\text{Wage}\times\text{Hours of work} \]

OpenStudy (anonymous):

? =5.25*1 ??

OpenStudy (anonymous):

m=5.25h..so it would be m=5.25(2)+15..so its $25.50..